x+y=4 Complete these ordered pairs for this equation
(0, ) (2, ) ( ,0)

Answer by nerdybill(7384)

x+y=4 Complete these ordered pairs for this equation
For each of the ordered pairs (x,y) they give you either the x or y-value. Substitute into the equation to find the missing value:
For: (0, )
x+y=4
0+y=4
y = 4
(0,4)
.
.
For: (2, )
x+y=4
2+y=4
y = 2
(2, 2)
.
.
For: ( ,0)
x+y=4
x+0=4
x = 4
(4 ,0)